OpenStudy (imamermaid29):

PLEASE HELP ME! find the value for x that makes the equation -3(x +10) +2x = 75 true.

OpenStudy (anonymous):

-3(x+10)+2x=75 -3x-30+2x=75 -x-30=75 -x=75+30 x=(-105)

OpenStudy (anonymous):

my medal

OpenStudy (imamermaid29):

Thank you so much !! :) @ambujverma